# Lightspeed Data Compute for the Space Era

### v1

@@ III System Overview
Figure 1: SpaceCoMP Architecture
@@ III-A Request Flow
A ground station initiates processing by submitting a job request to the nearest visible satellite, termed the Line-of-Sight (LOS) node (Figure 1). The request specifies:
@@
  • A Collect task: processes sensor configuration and data acquisition parameters.
  • A Map task: processes local-compute logic applied to each collected data partition.
  • A Reduce task: executes aggregation logic across all map outputs before a result downlink.
@@ III-B Processing Phases
Phase 1: Collect. The coordinator identifies satellites whose ground projections intersect the AOI. These collector nodes are instructed to acquire sensor data (e.g., capture imagery, record measurements). The data remains local to each collector, avoiding immediate network transfer.
Phase 2: Map. A subset of satellites within or near the AOI are designated as mapper nodes. The coordinator solves a bipartite matching problem (Section IV-B) to assign each collector’s data to a mapper, minimizing total transfer cost. Mappers execute the user-defined map function on their assigned data partitions, producing intermediate results.
Phase 3: Reduce. A single reducer node, selected to minimize aggregate transfer distance from all mappers (Section IV-B3), collects intermediate results and applies the reduce function. The final output is routed to the LOS node for downlink to the ground station.
@@ III-C Design Rationale
  • Topology awareness: Unlike terrestrial MapReduce, where data center topology is static, SpaceCoMP’s scheduler accounts for time-varying inter-plane distances when computing placement costs.
  • Downlink optimization: The reduce phase can compress petabytes of raw observations into megabytes of actionable intelligence before the bandwidth-constrained RF downlink.
@@ VI Discussion and Conclusion
We also discussed the architecture and baseline protocol, as well as ran simulations to validate the model. We also contributed a novel routing protocol, a map task allocation scheduler, and a reduce placement strategy.

### v2

@@ III System Overview
Figure 1: SpaceCoMP Architecture. Collectors observe the AOI via onboard sensors (wavy lines), exchange data with mappers and reducer via optical ISLs (solid lines), and deliver results to ground via RF links (dotted lines).
@@ III-A Request Flow
A ground station initiates processing by submitting a job request to the nearest visible satellite, termed the Line-of-Sight (LOS) node. The request specifies:
@@
  • A Collect task: sensor configuration and data acquisition parameters.
  • A Map task: local-compute logic applied to each collected data partition.
  • A Reduce task: aggregation logic combining all map outputs before downlink.
@@ III-B Processing Phases
Phase 1: Collect. The coordinator identifies satellites whose ground projections intersect the AOI. These collector nodes observe the AOI using onboard sensors (e.g., optical cameras, SAR, spectrometers) and store the acquired data locally, avoiding immediate network transfer.
Phase 2: Map. A subset of satellites within or near the AOI are designated as mapper nodes. The coordinator solves a bipartite matching problem (Section IV-B) to assign each collector’s data to a mapper, minimizing total transfer cost over ISL links. Mappers execute the user-defined map function on their assigned data partitions, producing intermediate results.
Phase 3: Reduce. A single reducer node, selected to minimize aggregate transfer distance from all mappers (Section IV-B3), collects intermediate results via ISL and applies the reduce function. The final output is routed to the LOS node for RF downlink to the ground station.
@@ III-C Design Rationale
  • Topology awareness: Unlike terrestrial MapReduce where datacenter topology is static, SpaceCoMP’s scheduler accounts for time-varying inter-plane distances when computing placement costs.
  • Downlink optimization: The reduce phase compresses terabytes of raw observations into megabytes of actionable intelligence before the bandwidth-constrained RF downlink.
@@ VI Discussion and Conclusion
Furthermore, we discussed the architecture and baseline protocol, as well as ran simulations to validate the model. Additionally, we contributed a novel routing protocol, a map task allocation scheduler, and a reduce placement strategy.
@@
Acknowledgments
This work was supported and funded by the Swedish National Space Agency (SNSA) as part of the LeoDOS project (https://www.leodos.org) under contract DNR 2025-00306.